# Image deconvolution and PSF reconstruction with STARRED: a wavelet-based two-channel method optimized for light curve extraction

### v1

@@ Abstract
We present starred, a Point Spread Function (PSF) reconstruction, two-channel deconvolution, and light curve extraction method designed for high-precision photometric measurements in imaging time series. An improved resolution of the data is targeted rather than an infinite one, thereby minimizing deconvolution artifacts. In addition, starred performs a joint deconvolution of all available data, accounting for epoch-to-epoch variations of the PSF and decomposing the resulting deconvolved image into a point source and an extended source channel. The output is a deep sharp frame combining all data, and the photometry of all point sources in the field of view as a function of time. Of note, starred also provides exquisite PSF models for each data frame. We showcase three applications of starred in the context of the imminent LSST survey and of JWST imaging: i) the extraction of supernovae light curves and the scene representation of their host galaxy, ii) the extraction of lensed quasar light curves for time-delay cosmography, and iii) the measurement of the spectral energy distribution of globular clusters in the “Sparkler”, a galaxy at redshift z=1.378 strongly lensed by the galaxy cluster SMACS J0723.3-7327. starred is implemented in jax, leveraging automatic differentiation and GPU acceleration. This enables rapid processing of large time-domain datasets, positioning the method as a powerful tool for extracting light curves from the multitude of lensed or unlensed variable and transient objects in the Rubin-LSST data, even when blended with intervening objects.
@@ 1 Introduction
Wavelet regularization introduces a minimally-informative prior to penalize the presence of non-zero elements in the wavelet domain, and is therefore connected with sparsity. An isotropic wavelet basis in which astronomical objects are well represented is the Starlets (Starck et al., 2015), meaning that they can be expressed as a linear combination of a few atoms of the Starlet dictionary. They also facilitate image denoising, since the different scales resulting from the Starlet transform can be penalized differently, with the noise being mostly represented in the finer wavelet scale.
In this paper, we propose to use starred (Michalewicz et al., 2023) to perform both the PSF reconstruction and the image deconvolution steps, based on the principles of MCS/Firedec, but with various key enhancements. These include the use of a family of isotropic wavelets (Starlets) specifically designed to represent astronomical objects, automatic differentiation, accelerated execution, scalability on Graphics Processing Units (GPUs), and the capability to handle multi-epoch deconvolution.
The method is very general and has a wide range of applications. These include but are not limited to light curve extraction of variable objects such as Active Galactic Nuclei (AGN) or supernovae, point source photometry in crowded fields, or the quasar/host decomposition problem (see e.g., Leist et al., 2023, for recent comparison between existing methods). Here, we select three science cases to demonstrate the capabilities of starred: i) light curve extraction and scene modeling of Type Ia Supernovae, ii) time-delay measurement in lensed quasars and iii) Spectral Energy Distribution (SED) measurement of high-redshift globular clusters. We apply this method to simulated Rubin-LSST time series and to real JWST observations, assessing the performance of the algorithm in terms of photometric and astrometric accuracy, and computational speed.
@@ 2.4 JAX implementation
Unlike deconvolution, PSF reconstruction cannot be performed for all the data at once and needs to be performed sequentially, as the PSF is expected to vary from epoch to epoch. However, this is easily parallelizable. Reconstructing the PSF for 6 $\times$ 32$\times$ 32 star cutouts takes approximately 5 seconds on a GPU, to which one should add 5 seconds of compilation at first initialization. If the method is deployed on a large scale, the processing speed may not be sufficient for handling millions of cutouts across thousands of epochs in the LSST survey. The main bottleneck in this process is not so much the optimization runtime, which fully leverages GPU capabilities, but rather the input/output (I/O) and pre-processing times, which are CPU-dependent. However, the I/O runtime and the CPU/GPU memory transfer could be largely optimized on a high-performance computing cluster. Moreover, if computational speed remains a concern, a feasible alternative would be to use the pre-calculated PSF models that are provided as part of the LSST image data products, reserving the PSF reconstruction method solely for the highest priority light curve extractions. However, this may come with a loss in photometric accuracy; further investigation is needed but this might just be a necessary trade-off for certain science cases.
@@ 3.1 Metric definitions
[…]
@@ 3.2 Results
In terms of PSF fidelity, we obtain a $\mathrm{MAE}_{\rm PSF}$ of $3.80\times 10^{-6}$, $2.43\times 10^{-5}$ and $1.70\times 10^{-5}$ for starred, psfr and photutils respectively. This corresponds to an improvement factor of 6.4 and 4.5 in comparison to psfr and photutils, respectively.
@@ 4.1 Example 1: scene modeling of supernova light curves
[…]
@@
[…]
@@ Results:
We confirm the point source nature of 13 out of the 18 compact objects modeled in the point source channel. Source ID 5, 6, 7, 11 and 12 are probably extended as some flux residuals are visible in the extended channel near the point sources. From the deblended photometry, we show the SED of these objects in Figure 10, while Table 3 collects all the measurements. The aperture photometry measured in Mowla et al. (2022) is shown with dotted lines in Figure 10. As our photometric measurements mostly agree for the unblended systems, we observe important differences for the source ID 4, 5, 7, 8, 10, 11, 12. The difference for the most blended system ID 7, 11 and 12 located in the center of the galaxy reach a factor of 2 to 3 difference, highlighting the need of properly modeling the background emission to obtain accurate photometry. Source ID 5 and 6 are extended, which explains the difference between the PSF photometry performed by starred and the aperture photometry of Mowla et al. (2022). Some of the newly measured point sources (ID 13 to 18) have similar SEDs than the globular clusters identified in (Mowla et al., 2022) (ID 1, 2, 4, 8 and 10). Future NIRSpec IFU data will confirm (or not) our findings.
@@ 6 Conclusion
    It introduces a new empirical PSF reconstruction technique, also based on sparse regularization. starred outperforms other empirical PSF reconstruction algorithms such as photutils or psfr in terms of the fidelity of the reconstructed PSF, photometric accuracy and astrometric precision. This is due to the sparse regularization, which effectively denoises the reconstructed PSF. starred also leverages the small subpixel shifts between stars relative to the pixel grid to reconstruct a subsampled version of the PSF. Obtaining a subsampled PSF is fundamental given the principle of the method aiming at an improved resolution rather than an infinite one, which requires that the deconvolved model, and therefore the PSF, can be measured on a fine grid of pixel. This is especially true with poorly sampled space data (see e.g., Symons et al., 2021).
As implemented, the method has a wide range of applications as it uses a generic PSF reconstruction algorithm and a robust two-channel deconvolution method. It is optimized for light curve extraction but of course also works with single-epoch observations. In this study we limit ourselves to three main science cases using simulated LSST images and real JWST observations, arriving to the following findings:
@@ Appendix C Regularization of the point source channel
The authors would like to thank Aymeric Galan for his precious advice on the development of starred and the CANUCS team for providing us access to their data. MM acknowledges support by the SNSF (Swiss National Science Foundation) through mobility grant P500PT_203114. KM acknowledges support from the President’s PhD Scholarship at Imperial College London. FC acknowledges support from SNSF.

### v2

# Image deconvolution and PSF reconstruction with STARRED: a wavelet-based two-channel method optimized for light-curve extraction

@@ Abstract
We present starred, a Point Spread Function (PSF) reconstruction, two-channel deconvolution, and light curve extraction method designed for high-precision photometric measurements in imaging time series. An improved resolution of the data is targeted rather than an infinite one, thereby minimizing deconvolution artifacts. In addition, starred performs a joint deconvolution of all available data, accounting for epoch-to-epoch variations of the PSF and decomposing the resulting deconvolved image into a point source and an extended source channel. The output is a high signal-to-noise, high resolution frame combining all data, and the photometry of all point sources in the field of view as a function of time. Of note, starred also provides exquisite PSF models for each data frame. We showcase three applications of starred in the context of the imminent LSST survey and of JWST imaging: i) the extraction of supernovae light curves and the scene representation of their host galaxy, ii) the extraction of lensed quasar light curves for time-delay cosmography, and iii) the measurement of the spectral energy distribution of globular clusters in the “Sparkler”, a galaxy at redshift z=1.378 strongly lensed by the galaxy cluster SMACS J0723.3-7327. starred is implemented in jax, leveraging automatic differentiation and GPU acceleration. This enables rapid processing of large time-domain datasets, positioning the method as a powerful tool for extracting light curves from the multitude of lensed or unlensed variable and transient objects in the Rubin-LSST data, even when blended with intervening objects.
@@ 1 Introduction
Wavelet regularization introduces a minimally-informative prior to penalize the presence of non-zero elements in the wavelet domain. An isotropic wavelet basis in which astronomical objects are well represented, i.e. sparsely represented, is the starlets (Starck et al., 2015), meaning that they can be expressed as a linear combination of a few atoms from the starlet dictionary. A sparse representation of the signal is a desired property as it facilitates image denoising. Since the noise is not sparse in the starlet domain, it can be reduced by reconstructing the image only from the most significant coefficients, in a technique called image thresholding. The different scales resulting from the starlet transform can be penalized differently, with the noise being mostly represented in the finer wavelet scale.
In this paper, we present the mathematical framework as well as several tests and science applications of starred, a method we previously introduced in Michalewicz et al. (2023). starred performs both the PSF reconstruction and the image deconvolution steps, based on the principles of MCS (Magain et al., 1998) and Firedec (Cantale et al., 2016), but with various key enhancements. These include the use of a family of isotropic wavelets (starlets) specifically designed to represent astronomical objects, automatic differentiation, accelerated execution, scalability on Graphics Processing Units (GPUs), and the capability to handle multi-epoch deconvolution.
The method is very general and has a wide range of applications. These include but are not limited to light curve extraction of variable objects such as Active Galactic Nuclei (AGN) or supernovae, point source photometry in crowded fields, or the quasar/host decomposition problem (see e.g., Leist et al., 2024, for recent comparison between existing methods). Here, we select three science cases to demonstrate the capabilities of starred: i) light curve extraction and scene modeling of Type Ia Supernovae, ii) time-delay measurement in lensed quasars and iii) Spectral Energy Distribution (SED) measurement of high-redshift globular clusters. We apply this method to simulated Rubin-LSST time series and to real JWST observations, assessing the performance of the algorithm in terms of photometric and astrometric accuracy, and computational speed.
@@ 2.4 JAX implementation
Unlike deconvolution, PSF reconstruction cannot be performed for all the data at once and needs to be performed sequentially, as the PSF is expected to vary from epoch to epoch. However, this is easily parallelizable. Reconstructing the PSF for 6 $\times$ 32$\times$ 32 star cutouts takes approximately 2.5 seconds on a GPU. If the method is deployed on a large scale, the processing speed may not be sufficient for handling millions of cutouts across thousands of epochs in the LSST survey. The main bottleneck in this process is not so much the optimization runtime, which fully leverages GPU capabilities, but rather the input/output (I/O) and pre-processing times, which are CPU-dependent. However, the I/O runtime and the CPU/GPU memory transfer could be largely optimized on a high-performance computing cluster. Moreover, if computational speed remains a concern, a feasible alternative would be to use the pre-calculated PSF models that are provided as part of the LSST image data products, reserving the PSF reconstruction method solely for the highest priority light curve extractions. However, this may come with a loss in photometric accuracy, if the PIFF PSF modeling approach that is planned for the LSST image processing pipeline follows that of the Dark Energy Survey (Jarvis et al., 2021). Further investigation is needed but this might just be a necessary trade-off for certain science cases.
@@ 3.1 Metric definitions
where $\boldsymbol{\hat{c}}_{k}$ is the estimated astrometric position of the kth star and $\boldsymbol{c}_{k}$ is its input position in the simulation. $\boldsymbol{b}$ is the astrometric bias defined as:
@@ 3.2 Results
In terms of PSF fidelity, we obtain a $\mathrm{MAE}_{\rm PSF}$ of 3.75×$\times 10^{-6}$, $1.25\times 10^{-5}$ and $1.70\times 10^{-5}$ for starred, psfr and photutils respectively. This corresponds to an improvement factor of 3.3 and 4.5 in comparison to psfr and photutils.
@@ 4.1 Example 1: scene modeling of supernova light curves
After the 30th epoch, we inject a varying point source in the galaxy to simulate a SNIa explosion. The light curves are interpolated from piscola (Müller-Bravo et al., 2022) standard SNIa template and scaled to reach a peak r-band magnitude of 19. We test 4 different locations of the SNIa in the galaxy, marked as white dots on the bottom right panel of Figure 5, to investigate the effect of the local surface brightness on the deblending performance.
@@
Location SB [mag.arsec-2] $P_{\rm starred}$ [mag] $P_{\rm DIA}$ [mag] $A_{\rm starred}$ [mag] $A_{\rm DIA}$ [mag] $\mathrm{P}_{\rm astro}$ [pixel]
@@ Results:
We confirm the point source nature of 13 out of the 18 compact objects modeled in the point source channel. Source ID 5, 6, 7, 11 and 12 are probably extended as some flux residuals are visible in the extended channel near the point sources. From the deblended photometry, we show the SED of these objects in Figure 10, while Table 3 collects all the measurements. The aperture photometry measured in Mowla et al. (2022) is shown with dotted lines in Figure 10. As our photometric measurements mostly agree for the unblended systems, we observe important differences for the source ID 4, 5, 7, 8, 10, 11, 12. The difference for the most blended system ID 7, 11 and 12 located in the center of the galaxy reach a factor of 2 to 3 difference, highlighting the need of properly modeling the background emission to obtain accurate photometry. Source ID 5 and 6 are extended, which explains the difference between the PSF photometry performed by starred and the aperture photometry of Mowla et al. (2022). For the other point sources, we are able to deblend them from the extended emission for the following two reasons. First, the degeneracy between the extended background and the point sources is not perfect, as the shape of the PSF is known to high precision. Therefore, all the features in the PSF (including the diffraction spikes, the Airy rings, etc…) are used to constrain the amplitude of the point source. Second, the starlet prior is informing the extended source morphology “under” the point sources. PSF features wrongly attributed to the extended source channel are penalized by this prior. Some of the newly measured point sources (ID 13 to 18) have similar SEDs than the globular clusters identified in (Mowla et al., 2022) (ID 1, 2, 4, 8 and 10). Future NIRSpec IFU data will confirm (or not) our findings.
@@ 6 Conclusion
    It introduces a new empirical PSF reconstruction technique, also based on sparse regularization. starred outperforms other empirical PSF reconstruction algorithms such as photutils or psfr in terms of the fidelity of the reconstructed PSF, photometric accuracy and astrometric precision. This is due to the sparse regularization, which effectively denoises the reconstructed PSF. starred leverages the small subpixel shifts between stars relative to the pixel grid to reconstruct a subsampled version of the PSF. Obtaining a subsampled PSF is fundamental given the principle of the method aiming at an improved resolution rather than an infinite one, which requires that the deconvolved model, and therefore the PSF, can be measured on a fine grid of pixel. This is especially true with poorly sampled space data (see e.g., Symons et al., 2021). We also emphasize that the high-quality PSF reconstruction that STARRED offers has diverse science applications, including lens modelling (Galan et al., 2024) or shear measurements (Finner et al., 2023).
As implemented, the method has a wide range of applications as it uses a generic PSF reconstruction algorithm and a robust two-channel deconvolution method. It is optimized for light curve extraction of point-like objects but of course also works with single-epoch observations or without any point sources in image, still providing an optimal way of combining multi-epoch observations with variable PSF. In this study we limit ourselves to three main science cases using simulated LSST images and real JWST observations, arriving to the following findings:
@@ Appendix C Regularization of the point source channel
The authors would like to thank Aymeric Galan for his precious advice on the development of starred and the CANUCS team for providing us access to their data. MM acknowledges support by the SNSF (Swiss National Science Foundation) through mobility grant P500PT_203114. KM acknowledges support from the President’s PhD Scholarship at Imperial College London. FC acknowledges support from SNSF. The JWST data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/jgpd-kj32 (catalog DOI: 10.17909/jgpd-kj32)